OpenStudy (anonymous):

One number is 7 less than another. The product of the numbers is 44. Find the numbers.

OpenStudy (campbell_st):

well let x be one number, so the other number is x - 7 so x(x-7) = 44 which is \[x^2 - 7x - 44 = 0\] just solve the quadratic for the solutions

OpenStudy (anonymous):

x(x-7)=44 x^2-7x-44=0 (x-11)(x+4)=0 I got X=11, -4 but my teacher said there was 4 answers

OpenStudy (campbell_st):

well thats not strictly correct as the 2nd number is dependent on the 1st... so there are 2 unique 1st numbers.... and the 2nd number depends on the 1st. the other method let x be the smaller number... so the larger number is x + 7 so x(x + 7) = 44 or \[x^2 + 7x - 44 = 0\]

OpenStudy (anonymous):

Would they be (11,4) and (-4,11)?

OpenStudy (campbell_st):

so the solutions are 11 and 4 -4 and - 11

OpenStudy (anonymous):

Gratzi!

OpenStudy (anonymous):

Would there always be 2 solutions to this type of problem or just depending on the numbers?

OpenStudy (campbell_st):

well a quadratic always has 2 solutions... they may be equal... as in a perfect square... or complex... or real... but always 2

OpenStudy (anonymous):

mkay, thanks.